OpenStudy (anonymous):

please help!!!!!!! let letters of the alphabet be A = 01, B = 02, .. Z = 26. For RSA encryption let n = 119 and e = 5. Encrypt letter W. Show work. Thanks!!

myininaya (myininaya):

So the message is W And we want to encrypt this message Do you know the rsa encryption function?

OpenStudy (anonymous):

f(p) = (p+k) (mod 26)

myininaya (myininaya):

\[E(message=m)=m^e (\mod n)\]

myininaya (myininaya):

http://en.wikipedia.org/wiki/RSA_(algorithm)

myininaya (myininaya):

\[E(m)=m^e \mod n\] We have n=119 and e =5 So we have \[E(m)=m^5 \mod 119\]

myininaya (myininaya):

Now I believe W is the 23rd number in the alphabet So W=23 Replace m with 23 and evaluate to get the ciphertext

myininaya (myininaya):

And that is it :)

myininaya (myininaya):

According to above you said W is 23

myininaya (myininaya):

" A = 01, B = 02, .. Z = 26"

myininaya (myininaya):

A=01 B=02 C=03 D=04 E=05 F=06 G=07 and so

OpenStudy (anonymous):

so the answer would have to be a value between 0 and 26?

OpenStudy (anonymous):

sorry 1 and 26?

myininaya (myininaya):

Should be a value between 0 and n

OpenStudy (anonymous):

oh okay that makes sense, i got 109. Thanks!

myininaya (myininaya):

Ok great! :)

OpenStudy (anonymous):

i just thought the encryption might have to be an alphabet value, so its ok if its not?

myininaya (myininaya):

I believe 109 is right :)